OpenStudy (czesc):

INSTANT FAN AND MEDAL EASY For what value of x is the slope of the curve undefined for the graph defined by x = 10 - t^2, y = t^3 - 12t?

OpenStudy (solomonzelman):

\(\color{#000000}{ \displaystyle y=t^3-12t }\) \(\color{#000000}{ \displaystyle x=-t^2+10 }\) \(\color{#000000}{ \displaystyle \frac{dy}{dx}=\frac{\displaystyle \frac{dy}{dt}}{\displaystyle \frac{dx}{dt}} }\)

OpenStudy (solomonzelman):

Find, \(\color{#000000}{ \displaystyle y' }\) and \(\color{#000000}{ \displaystyle x' }\).

OpenStudy (czesc):

@SolomonZelman thanks for helping, what should I do next?